OpenStudy (anonymous):

Sonja's house is 4 blocks west and 1 block south of the center of town. Her school is 3 blocks east and 2 blocks north of the center of town. What is the direct distance from Sonja's house to her school? (Hint: The center of town should be the origin, and north is up.)

OpenStudy (paxpolaris):

 Find the distance between (-4,-1) & (3,2)

OpenStudy (paxpolaris):

\[\Large = \sqrt{\left[ 4-(-3) \right]^2+\left[ 2-(-1) \right]^2}=\sqrt{7^2+3^2}=\sqrt{58}\]